# Superhuman AI Disclosure: Impacts on Toxicity, Fairness, and Trust Vary by Expertise and Persona Attributes

### v1

@@ Abstract.
As artificial intelligence demonstrates surpassing human performance across real-world tasks, disclosing “superhuman” capabilities poses challenges for fairness, accountability, and trust. To investigate how transparency impacts attitudes and perceptions, we introduce a grounded and validated set of synthetic personas reflecting diverse fairness concerns and technology acceptance levels. Then we evaluate responses in two contrasting domains: (1) a competitive player in StarCraft II, where strategy and high-skill gameplay often elicit toxic interactions, and (2) a cooperative personal-assistant in providing information. Across numerous interactions spanning persona profiles, we test non-disclosure versus explicit “superhuman” labeling under controlled game outcomes and usage contexts. Our findings reveal sharp domain-specific effects: in StarCraft II, explicitly labeling AI as “superhuman,” novice personas who learned of it reported lower toxicity and higher fairness—attributing defeat to advanced skill rather than hidden cheating—whereas expert personas found the disclosure statements irksome but still less deceptive than non-disclosure. Conversely, in the LLM as personal-assistant setting, disclosure of superhuman capabilities improved perceived trustworthiness, though it risked AI overreliance among certain persona segments. We release Dataset X-containing persona cards—including profile attributes, disclosure prompts, and detailed interaction logs, accompanied by reproducible protocols and disclaimers for adapting them to diverse tasks. Our results demonstrate that transparency is not a cure-all: while it reduces suspicion and enhances trust in cooperative contexts, it may inflame resistance or disappointment in competitive domains. Our work clarifies that while explicit disclosure can counter hidden advantages and reinforce accountability, its success hinges on careful tailoring to user expectations, domain norms, and fairness objectives.
Model Behavior, Transparency, Disclosure, LLMs, Game AI, StarCraft II, Contextual Bias
@@ 1. INTRODUCTION
AI systems increasingly operate behind the scenes in domains such as clinical decision support, e-commerce, and multiplayer gaming, prompting critical questions about fairness, accountability, and transparency (Friedler et al., 2021; Pagano et al., 2023). A striking example is DeepMind’s AlphaStar, which achieves top-level play in the real-time strategy (RTS) game StarCraft II via inhumanly fast maneuvers and near-perfect resource management (Silver et al., 2017; Vinyals et al., 2019). While these achievements highlight significant AI capabilities, they also raise concerns about how humans perceive and place trust in AI with a seemingly insurmountable machine advantage.
@@
In StarCraft II and other multiplayer gaming contexts, much research has addressed overt harassment and toxic gaming behavior (Laato et al., 2024; Kou, 2020), yet there is insufficient understanding of how subtler manifestations of AI dominance affect user trust and fairness perceptions—particularly when systems like AlphaStar deploy superhuman capabilities (Vinyals et al., 2019; Justesen and Risi, 2017). This gap becomes more critical as transparency about AI advantages can simultaneously mitigate suspicion or provoke frustration (Justesen and Risi, 2017; Eslami et al., 2016), depending on whether users see these capabilities as rightful skill or an unfair edge. This issue extends beyond competitive gaming, impacting any scenario where humans interact with advanced AI. The imperative for transparency in AI transcends adversarial contexts and encompasses scenarios where AI functions as an ally. Even in cooperative settings, the application of advanced AI presents challenges. Empirical evidence from customer service suggests that explicit disclosure of AI utilization evoke perceptions of manipulation (Seeger et al., 2021; Cai et al., 2019; Gillespie, 2018; van der Goot et al., 2024; Adam et al., 2021). In recommendation systems, despite its intended helpfulness, highly accurate predictions and personalized outputs may be construed as a violation of privacy (Yang et al., 2023; Zhu and Grover, 2022; Logg et al., 2019). Despite ongoing efforts to improve AI transparency, it remains unclear whether disclosing “superhuman” AI may either foster trust calibration or confirm doubts about cheating and undue influence (Holstein et al., 2022). Even with carefully framed disclosures, transparency alone does not guarantee optimal trust calibration (Jiang, 2024; Zhang et al., 2024; Hagendorff and Wezel, 2023). AI explanations confuse users if explanations exceed their expertise  (Bansal et al., 2021); and Novice-Intermediate might overly trust a system perceived as infallible leading to AI overreliance  (Kocielnik et al., 2019; Schmitt et al., 2024). Consequently, a deeper investigation into how disclosure interacts with user expertise and domain context is critical for achieving balanced and ethical AI deployment.
@@
To address the laid out challenges posed by disclosing ”superhuman” AI capabilities, our research investigates the following question: RQ1: How does concealment versus disclosing superhuman AI capabilities influence user perceptions of fairness, toxicity, and trust? RQ2: And to what extent do user characteristics attributes to personality or notions of play style, particularly expertise level, moderate the impact of superhuman AI disclosure on user attitudes and behaviors?
@@
Naturally suited for investigating our research questions is StarCraft II, a real-time strategy (RTS) game that provides a high-skill, adversarial, human-computer interaction environment that reliably elicits intense strategic interactions and strong emotional investment from players. Superhuman capabilities are present in this setting, as participants face an opponent with inhumanly high actions per minute (APM) and strategic choice for early dominance. We model responses from casual-to-mid-level and expert synthetic persona players, varying whether they receive disclosure about the AI’s superhuman capabilities in order to answer our first research question. As for our second research question, each synthetic persona incorporates defined cognitive architectures and belief-updating mechanisms, allowing us to examine how attributes such as certain parts of the population with varying openness and degrees of socio-technical acceptance, and prior assumptions, shape reactions to superior AI performance. In order to further back up our data collection, we analyze quantitative telemetry (APM logs, resource utilization) and post-match interviews to ground the synthetic data generation and gauge fairness, enjoyment, and frustration. We tracked user reactions and perceptions through persona-based data generation, employing detailed cognitive architecture attributes. We then analyzed post-task interview responses to ground the synthetic rating data generation.
Our results show that disclosure powerfully shapes perceptions, but its effects diverge by user expertise and domain. In StarCraft II, expert personas who learned of the AI’s superhuman capacity reframed formidable gameplay as legitimate skill, significantly reducing reported toxicity (M=5.00 down to M=2.70). By contrast, novice players sometimes grew resentful or over-reliant, expecting flawless performance. This sparked curiosity about establishing baselines for such results, as there might be differences between opponent and ally interactions. Therefore we also ran the same experimental steps with a more familiar conversational assistant powered by a large language model (LLM), which provided information prompts in the Appendix. In the LLM scenario, experts became more critical upon discovering the AI’s advanced status, whereas Novice-Intermediate displayed heightened trust. These differences reveal that while transparency can dispel suspicions in adversarial settings, it also increases scrutiny in collaborative tasks—a context-dependent dynamic that calls for nuanced disclosure strategies.
@@
Alongside these findings, we introduce Dataset X, a corpus of systematically generated persona interactions grounded in well-documented cognitive architectures. This resource offers high reproducibility and fine-grained control of factors like skill, domain norms, and disclosure levels. By analyzing how superhuman advantages influence fairness, toxicity, and trust, Dataset X addresses both the ethical challenges and the methodological demands of studying advanced AI across different tasks without confounding biases. Ultimately, our work illuminates the delicate interplay between transparency and user expectations and underscores the need for adaptive design strategies that consider the user’s domain context and level of expertise.
@@ 2. RELATED WORK
2.1. Perceived Unfairness and Toxicity in Competitive Games
@@
Fairness and toxicity have been extensively studied in human-to-human settings, focusing on verbal harassment, abuse, or trolling (Kuznekoff and Rose, 2013; Kou, 2020; Laato et al., 2024). However, AI opponents introduce new dimensions of mechanical or strategic toxicity. In real-time strategy (RTS) games such as StarCraft II and Dota 2, AI systems frequently exceed professional human norms of around 300 actions per minute (APM) (Vinyals et al., 2019), prompting accusations of inhuman or unfair play (Peters et al., 2020; Justesen and Risi, 2017). This is consistent with research on learning from human feedback, where high APM in AI players was perceived as unfair (Justesen and Risi, 2017). Developers sometimes cap AI actions, as in limiting AlphaStar’s APM, to avoid overwhelming players (Vinyals et al., 2019; Berner et al., 2019), yet these caps rarely address whether explicitly stating ”this game AI is superhuman” shifts player perceptions of cheating (Pfau et al., 2024; Kim and Ortiz, 2024; Kocielnik et al., 2019). The potential for AI to exploit game mechanics, often referred to as ”reward hacking,” can further exacerbate perceptions of unfairness (Amodei et al., 2016; Krakovna et al., 2020). Reward hacking can also yield seemingly illegitimate strategies that exploit the game’s reward function (Amodei et al., 2016; Krakovna et al., 2020), compounding notions of unfair advantage. A previous study has shown that toxicity arises from multiple factors, including competitive mechanics and perceived misalignment with social norms (Beres et al., 2021). Our research extends this by investigating how labeling an AI as superhuman might exacerbate negativity if players interpret any deviation from human-like performance as cheating. These concerns parallel the alignment problem (Hadfield-Menell et al., 2017; Lehman et al., 2018), wherein AI systems pursue narrow objectives that may conflict with user expectations of fair or balanced gameplay. Henceforth, our work builds on these insights by examining how superhuman disclosure interacts with these established challenges in competitive gaming AI, and takes inspiration from work on how the relation to human competence relates to the reliance on AI (He et al., 2023).
2.2. Effects of Disclosure as Transparency on Trust
@@
Research on AI disclaimers spans a variety of interaction modes, often focusing on large language models (LLMs) that generate near-human text (Gillespie, 2018; Jiang, 2024). While disclosing “This is AI-generated” can help users calibrate trust, it does not consistently alter perceptions of toxicity (Zhang et al., 2024). Studies on blame attribution for algorithmic harm suggest that user reactions depend on factors such as perceived fairness, the harmfulness of the AI’s actions, and personal attitudes toward assigning blame to machines (Lima et al., 2023). Complicating matters, LLMs may self-moderate upon realizing they are under scrutiny, reflecting moral judgments embedded in their training data (Ganguli et al., 2022). Moreover, demonstrating trust on through complying with AI systems is not dictated solely by the accuracy or precision of the text output. For instance, anthropomorphic design cues have been shown to increase compliance with chat bot requests, yet explicit AI labeling can provoke skepticism or diminish persuasion if users feel misled (Adam et al., 2021). Related work on avatar disclosure indicates that revealing an avatar’s artificial nature can influence how intensely participants play or engage, especially when the avatar’s appearance suggests human likeness (Visser et al., 2024). However, in customer-service environments, disclosure of a chat bots artificial identity did not necessarily undermine social presence or reduce helpfulness, as users often prioritized perceived usefulness over identity status (van der Goot et al., 2024). User familiarity with AI also plays a significant role: individuals less experienced with AI may misjudge or overestimate toxicity in AI-generated content, while more experienced users are often unmoved by disclaimers (Hagendorff and Wezel, 2023). By contrast, some studies indicate that disclaimers can heighten user caution and prevent AI overreliance on the systems (Wischnewski et al., 2023; Jiang, 2024).
@@ 2.3. Conceptual Framework and Key Definitions
This scopes how we define and use the primary constructs examined in: toxicity, fairness, and trust. Although these terms appear throughout multiple literature, we adopt specific working definitions tailored to our experimental contexts. We define toxicity including toxic gaming behavior and provoking communication (biased or insulting responses)—whether automated or user-driven—that is perceived as harmful, abusive, or excessive and therefore annoying (Kou, 2020; Laato et al., 2024). In our StarCraft II scenario, toxicity may take the form of in-game chat or even gameplay actions that implies unfair superiority or demeans the opponent. Our measurement focuses on perceived negativity rather than explicit harassment alone. As for Fairness, building on established notions in social and algorithmic contexts (Selbst et al., 2019; Jacovi et al., 2021), we frame fairness as the extent to which an AI operates within recognized boundaries of normative behavior—abiding by rules, community standards, and user expectations. In StarCraft II, fairness specifically entails: (1) adherence to typical mechanical and strategic limitations, (2) avoidance of “unseen” or privileged information not accessible to a human opponent, and (3) reliance on tactics that conform to community norms (e.g., not exploiting known design loopholes). Recent work further highlights the necessity for AI systems to demonstrate a perceived absence of bias or undue manipulation in textual, visual, or interactive outputs (and actions, in our case). A key assumption in our study regarding fairness is that experienced players, given their higher socio-technical technology acceptance, may both recognize and tolerate more advanced AI capabilities—provided those capabilities remain within what they perceive as “fair play.” And for building on the concept of trust in human–AI interaction (Bansal et al., 2021) through two main indicators: the persona’s willingness to rely on the AI’s outputs (or play against it again, in the game scenario) and the degree of confidence that the AI is acting consistently with its disclosed capabilities. High trust emerges when the persona perceives the AI as reliable, transparent, and aligned with their expectations.

### v2

# Superhuman Game AI Disclosure: Expertise and Context Moderate Effects on Trust and Fairness

@@ Abstract.
As artificial intelligence surpasses human performance in select tasks, disclosing superhuman capabilities poses distinct challenges for fairness, accountability, and trust. However, the impact of such disclosures on diverse user attitudes and behaviors remains unclear, particularly concerning potential negative reactions like discouragement or overreliance. This paper investigates these effects by utilizing Persona Cards: a validated, standardized set of synthetic personas designed to simulate diverse user reactions and fairness perspectives. We conducted an ethics board-approved study (N=32), utilizing these personas to investigate how capability disclosure influenced behaviors with a superhuman game AI in competitive StarCraft II scenarios. Our results reveal transparency is double-edged: while disclosure could alleviate suspicion, it also provoked frustration and strategic defeatism among novices in cooperative scenarios, as well as overreliance in competitive contexts. Experienced and competitive players interpreted disclosure as confirmation of an unbeatable opponent, shifting to suboptimal goals. We release the Persona Cards Dataset, including profiles, prompts, interaction logs, and protocols, to foster reproducible research into human alignment AI design. This work demonstrates that transparency is not a cure-all; successfully leveraging disclosure to enhance trust and accountability requires careful tailoring to user characteristics, domain norms, and specific fairness objectives.
Transparency, Trust, Disclosure, Overreliance, Model Behavior, Contextual Bias, LLMs, Game AI, StarCraft II
@@ 1. INTRODUCTION
AI systems operate behind the scenes in domains such as clinical decision support, e-commerce, and multiplayer gaming, prompting critical questions about fairness, accountability, and transparency (Friedler et al., 2021; Pagano et al., 2023). An influential example is DeepMind’s AlphaStar, which achieves pro level play in the real-time strategy (RTS) game StarCraft II through inhumanly fast maneuvers and near-perfect resource management (Silver et al., 2017; Vinyals et al., 2019). Although these achievements highlight significant AI capabilities, they also raise concerns about how humans perceive and place trust in AI with a seemingly insurmountable machine advantage described as superhuman.
@@
In StarCraft II and other multiplayer gaming contexts, much research has addressed overt harassment and toxic gaming behavior (Laato et al., 2024; Kou, 2020), yet there is insufficient understanding of how subtler manifestations of AI dominance affect user trust and fairness perceptions particularly when systems like AlphaStar deploy superhuman capabilities (Vinyals et al., 2019). This gap becomes more critical as transparency about AI advantages can simultaneously mitigate suspicion or provoke frustration (Eslami et al., 2016), depending on whether users see these capabilities as rightful skill or an unfair edge. Transparency issues also arise in assistant scenarios where AI serves as an ally, as even in cooperative settings, the application of advanced AI presents challenges. Empirical evidence from customer service suggests that explicit disclosure of AI utilization evoke perceptions of manipulation (Seeger et al., 2021; Cai et al., 2019; Gillespie, 2018; van der Goot et al., 2024; Adam et al., 2021). In recommendation systems, despite its intended helpfulness, highly accurate predictions and personalized outputs may be construed as a violation of privacy (Yang et al., 2023; Zhu and Grover, 2022; Logg et al., 2019). Despite ongoing efforts to improve AI transparency, it remains unclear whether disclosing superhuman AI model capabilities may either foster trust calibration or confirm doubts about cheating and undue influence (Holstein et al., 2022). Even with carefully framed disclosures, transparency alone does not guarantee optimal trust calibration (Jiang, 2024; Zhang et al., 2024; Hagendorff and Wezel, 2023). AI explanations confuse users if explanations exceed their expertise  (Bansal et al., 2021); and novice players might overly trust a system perceived as infallible leading to AI overreliance  (Kocielnik et al., 2019; Schmitt et al., 2024b). Consequently, an investigation into how model capability disclosure affects user expertise and domain context is important for achieving ethical and responsible AI deployment. We do not yet understand how explicit disclosure of superhuman capabilities shapes user attitudes and behaviors across adversarial versus cooperative settings, for users with different expertise levels.
@@
To address the laid out challenges posed by disclosing superhuman AI capabilities, our research investigates the following question: RQ1: How does concealment versus disclosing superhuman AI capabilities influence user perceptions of fairness, toxicity, and trust? RQ2: And to what extent do user characteristics attributes to personality or notions of play style, particularly expertise level, moderate the impact of superhuman AI disclosure on user attitudes and behaviors?
Naturally suited for investigating our research questions is StarCraft II, a real-time strategy (RTS) game known to reliably elicit complex strategic behavior and emotional investment, making it particularly effective for exploring trust calibration and fairness perceptions under superhuman AI contexts. However, despite the carefully framed disclosures suggested by prior research, transparency alone may fail to consistently optimize trust calibration due to varying user expertise and predispositions toward technology. Our study hypothesizes that explicit disclosure of superhuman AI capabilities may heighten user frustration and suspicion, particularly among novices, whereas experts may respond with suboptimal strategy adaptations.
@@
Our research contribution to responsible AI involves developing a concept called Persona Cards, inspired by model cards. These cards serve as a tool to track user values and biases that influence ratings, enabling the creation of synthetic human participant data. The goal is to model responses from casual to mid-level and expert synthetic persona players, varying whether they are informed about the AI’s superhuman capabilities. This allows us to address our first research question: how do different levels of disclosure about the AI’s capabilities affect user reactions? As for our second research question, each synthetic persona incorporates defined cognitive architectures and belief updating mechanisms. This enables us to examine how various factors, such as different levels of openness and acceptance of technology among different segments of the population, and prior assumptions, shape reactions to superior AI performance. To further validate our data collection, we analyse quantitative telemetry (APM logs and resource utilisation) and post-match interviews to ground the synthetic data generation process and assess fairness, enjoyment, and frustration. We tracked user reactions and perceptions through persona-based data generation, incorporating detailed cognitive architecture attributes.
In order to address ecological validity and ethical considerations regarding synthetic persona data, we conducted an IRB approved experiment involving real human participants (N=32) playing against our pro-level StarCraft II game AI with variable APM rates. Participants were randomly assigned to either disclosure or non-disclosure conditions, ensuring that our hypothesis remained concealed while still informing them of the model’s advanced capabilities in the disclosure arm. Immediately after each match, we debriefed participants to mitigate potential frustration.
Our findings show that disclosure significantly shapes perceptions, but its effects vary on the user’s expertise and domain. In StarCraft II, novice players sometimes developed resentment or excessive reliance, expecting flawless performance. On the other hand, expert players who usually caution against superhuman play and experimented with novel strategies that could potentially outperform the AI, after being informed of the AI’s superhuman capabilities, concluded formidable gameplay as unbeatable. This highlights that humans are prone to be misled to pursue subpar goals, such as mini-games to prolong the game timer before defeat from the AI, instead of investing their time and effort in achieving their original and true objective to genuinely beat the AI. The results matched with the IRB approved study, as human participants also excessively relied on the system based on the researchers’ disclosure statement.
@@
Alongside findings, we introduce Persona Cards Dataset, a corpus of systematically generated persona interactions grounded in cognitive architectures, offering high reproducibility and controlled variations. Analyzing how superhuman advantages influence fairness, toxicity, and trust, Persona Cards Dataset addresses both the ethical challenges and the methodological demands of studying advanced AI across different tasks without confounding biases. Therefore, our study offers the following contributions to advance the responsible development and deployment of AI systems:
  • Methodological Contribution. We introduce Persona Cards, a novel systematic framework explicitly documenting user values and biases that inform rating behaviors, enabling more transparent, accountable, and fairness-aware synthetic persona generation for AI assessments.
  • Empirical Contribution. We empirically demonstrate that disclosure regarding an AI’s superhuman capabilities move fairness judgments validated across both adversarial and collaborative contexts.
  • Dataset Contribution. We contribute Persona Cards Dataset, an open source repository of persona card interactions grounded within defined cognitive architectures, enabling reproducible experimentation and cross-domain insights.
  • Practical Design Guidelines. We propose concrete adaptive design strategies, emphasizing balanced transparency tailored to user expertise, to practically foster equitable AI experiences and strengthen real-world accountability and fairness standards.
Section 2 reviews perceptions of fairness, toxicity in gaming, and disclosure as a mechanism for establishing trust. Section 3 details synthetic persona as a method for HCI research attribute manipulation. Section 4 presents the empirical findings, including statistical confirmation. Section 5 interprets these results for fairness and trust calibration. Section 6 concludes, and Section 7 reflects on ethics and limitations of our study.
@@ 2. RELATED WORK
2.1. Socio-Technical Governance and AI Policy
There is a movement of artificial intelligence technology that is beyond technical considerations towards robust socio-technical governance frameworks and effective AI policy. As AI systems become increasingly embedded in societal structures, there is a clear consensus that ensuring responsible development and deployment requires bridging the gap between high-level ethical principles and concrete, enforceable mechanisms. (Shneiderman, 2020) articulate this shift, arguing that abstract values like fairness, transparency, and accountability must be operationalized through specific laws, regulations.
A central theme in contemporary scholarship is the framing of AI governance as an inherently socio-technical challenge. This perspective in evaluating AI systems in isolation, instead considering the complex interplay between the technology, its developers, deploying organizations, end-users, and regulatory bodies (Selbst et al., 2019). Issues such as bias, safety, or trustworthiness are understood not merely as technical properties but as emergent outcomes of this broader ecosystem. (Manzini et al., 2024), for example, propose that establishing trust in advanced AI assistants necessitates evaluating evidence of competence and alignment across multiple levels: the AI system’s design features, the organizational practices of its creators (including ethics reviews and testing protocols), and the effectiveness of third-party governance mechanisms like external audits or certifications. This multi-level, contextual approach, examining AI within its intricate web of human and institutional interactions, is increasingly recognized as essential for effective governance (Selbst et al., 2019). Addressing the unique challenges posed by advanced AI systems, such as large language models and autonomous agents, has become a focal point for policy and research. Concerns regarding misinformation, market concentration, and potential risks have spurred the development of specific regulatory frameworks. Notable examples include the US NIST AI Risk Management Framework (AI, 2023) and the EU AI Act (Nikolinakos, 2023), both advocating for risk-based approaches that impose stricter requirements on high-impact applications.
2.2. Perceived Unfairness of Game AI Capability
@@
Fairness and toxicity have been extensively studied in human-to-human settings, focusing on verbal harassment, abuse, or trolling (Kuznekoff and Rose, 2013; Kou, 2020; Laato et al., 2024). However, AI opponents introduce new dimensions of mechanical or strategic toxicity. In real-time strategy (RTS) games such as StarCraft II and Dota 2, AI systems frequently exceed professional human norms of around 300 actions per minute (APM) (Vinyals et al., 2019), prompting accusations of inhuman or unfair play (Peters et al., 2020). This is consistent with research on learning from human pro play in StarCraft II, where high APM in AI players was perceived as unfair. Developers sometimes cap AI actions, as in limiting AlphaStar’s APM, to avoid overwhelming players (Vinyals et al., 2019; Berner et al., 2019), yet these caps rarely address whether explicitly stating ”this game AI is superhuman” shifts player perceptions of cheating (Pfau et al., 2024; Kim and Ortiz, 2024; Kocielnik et al., 2019). The potential for AI to exploit game mechanics, often referred to as reward hacking, can further exacerbate perceptions of unfairness (Amodei et al., 2016; Krakovna et al., 2020). Reward hacking can also yield seemingly illegitimate strategies that exploit the game’s reward function (Amodei et al., 2016; Krakovna et al., 2020), compounding notions of unfair advantage. A previous study has shown that toxicity arises from multiple factors, including competitive mechanics and perceived misalignment with social norms (Beres et al., 2021). Our research extends this by investigating how labeling an AI as superhuman might exacerbate negativity if players interpret any deviation from human-like performance as cheating. These concerns parallel the alignment problem (Hadfield-Menell et al., 2017; Lehman et al., 2018), wherein AI systems pursue narrow objectives that may conflict with user expectations of fair or balanced gameplay. Henceforth, our work builds on these insights by examining how superhuman disclosure interacts with these established challenges in competitive gaming AI, and takes inspiration from work on how the relation to human competence relates to the reliance on AI (He et al., 2023).
2.3. Effects of Disclosure on Trust
@@
Research on AI disclaimers spans a variety of interaction modes, often focusing on large language models (LLMs) (Gillespie, 2018; Jiang, 2024). While disclosing “This is AI generated” can help users calibrate trust, it does not consistently alter perceptions of toxicity (Zhang et al., 2024). Studies on blame attribution for algorithmic harm suggest that user reactions depend on factors such as perceived fairness, the harmfulness of the AI’s actions, and personal attitudes toward assigning blame to machines (Lima et al., 2023). Complicating matters, LLMs may self-moderate upon realizing they are under scrutiny, reflecting moral judgments embedded in their training data (Ganguli et al., 2022). Moreover, demonstrating trust on through complying with AI systems is not dictated solely by the accuracy or precision of the text output. For instance, anthropomorphic design cues have been shown to increase compliance with chat bot requests, yet explicit AI labeling can provoke skepticism or diminish persuasion if users feel misled (Adam et al., 2021). Related work on avatar disclosure indicates that revealing an avatar’s artificial nature can influence how intensely participants play or engage, especially when the avatar’s appearance suggests human likeness (Visser et al., 2024). However, in customer-service environments, disclosure of a chat bots artificial identity did not necessarily undermine social presence or reduce helpfulness, as users often prioritized perceived usefulness over identity status (van der Goot et al., 2024). User familiarity with AI also plays a significant role: individuals less experienced with AI may misjudge or overestimate toxicity in AI-generated content, while more experienced users are often unmoved by disclaimers (Hagendorff and Wezel, 2023). In contrast, other studies indicate that disclaimers can heighten user caution and prevent AI overreliance on the systems (Wischnewski et al., 2023; Jiang, 2024).
@@ 2.4. Conceptual Framework and Key Definitions
This scopes how we define and use the primary constructs examined in: toxicity, fairness, and trust. Although these terms appear throughout multiple literature, we adopt specific working definitions tailored to our experimental contexts. We define toxicity including toxic gaming behavior and provoking communication (biased or insulting responses) whether automated or user-driven that is perceived as harmful, abusive, or excessive and therefore annoying (Kou, 2020; Laato et al., 2024). In our StarCraft II scenario, toxicity may take the form of in-game chat or even gameplay actions that implies unfair superiority or demeans the opponent. Our measurement focuses on perceived negativity rather than explicit harassment alone. As for Fairness, building on established notions in social and algorithmic contexts (Selbst et al., 2019; Jacovi et al., 2021), we frame fairness as the extent to which an AI operates within recognized boundaries of normative behavior, abiding by rules, community standards, and user expectations. In StarCraft II, fairness specifically entails: (1) adherence to typical mechanical and strategic limitations, (2) avoidance of “unseen” or privileged information not accessible to a human opponent, and (3) reliance on tactics that conform to community norms (e.g., not exploiting known design loopholes). Recent work further highlights the necessity for AI systems to demonstrate a perceived absence of bias or undue manipulation in textual, visual, or interactive outputs (and actions, in our case).
A key assumption in our study regarding fairness is that experienced players, given their higher socio-technical technology acceptance, may both recognize and tolerate more advanced AI capabilities, provided those capabilities remain within what they perceive as “fair play.” And for building on the concept of trust in human–AI interaction (Bansal et al., 2021) through two main indicators: the persona’s willingness to rely on the AI’s outputs (or play against it again, in the game scenario) and the degree of confidence that the AI is acting consistently with its disclosed capabilities. High trust emerges when the persona perceives the AI as reliable, transparent, and aligned with their expectations.
Placement in the Literature
Prior research emphasizes the importance of operationalizing transparency, fairness, and accountability into enforceable socio-technical governance mechanisms and actionable policy frameworks (Shneiderman, 2020; Selbst et al., 2019; Nikolinakos, 2023; AI, 2023). Yet, existing approaches largely presume AI systems operate at or near human competence, leaving a notable gap when dealing explicitly with superhuman capabilities (Vaccaro et al., 2024; Schmitt et al., 2024a). This introduces novel complexities, particularly around fairness defined as adherence to community norms and ethical expectations (Jacovi et al., 2021) and trust calibration, aligning users’ expectations accurately with the system’s true capabilities (Lee and See, 2004; Bansal et al., 2021). The following Methods section details our empirical approach to answering these targeted research questions comprehensively.